# MalMixer: Few-Shot Malware Classification with Retrieval-Augmented Semi-Supervised Learning

### v2

@@ Appendix B Implementation Details
We use two Nvidia A30 GPUs during model training and evaluation. We set the random seed as 17 in all processes for reproducibility. We note that training and evaluation across all experimental settings required less than one hour each using BODMAS-20 with $1\%$ labeled data. In our experiments, we standardize all features in the BODMAS dataset through z-score scaling as in standard practice. For the encoder-decoder architecture, encoder $\phi_{i}$ comprises three layers, with each layer outputting 512 features. The corresponding decoder $\psi_{i}$ consists of three layers, mirroring the structure of the encoder in reverse order. Encoder $\phi_{n}$ has five layers, outputting 1024, 1024, 512, 512, and 512 features respectively. The corresponding the decoder $\psi_{n}$ reflects the encoder’s architecture in reverse order. Consequently, the embeddings $h_{u}^{d},h_{u}^{s},h_{i}^{d},h_{i}^{s}$ created by the encoder-decoder architecture are all represented as 256-dimensional vectors. For our retrieval-based augmentation, we use Faiss [59], an open-source library designed for scalable similarity search and clustering, to construct the k-nearest-neighbor graph and perform similarity searches with L2 distance. Top 5 neighbors are selected for the $k$-NN graph in the original feature space $\mathcal{S}$ and 5 candidates are selected for alignment-based malware selection. For the semi-supervised framework, unless otherwise noted, we use ResNet-12 [57] as the base classifier for MalMixer and all other semi-supervised methods evaluated. We empirically found that removing all BatchNorm layers in ResNet-12 leads to improved performance of MalMixer. For all training and evaluation of MalMixer, we set $\lambda=1$ in Equation (5) and linearly ramp up $\lambda$ to its maximum value [60]. We perform up-sampling through augmentation to each malware family in the labeled set to create a balanced $\mathcal{X}^{\prime}$. Each evaluation task is performed five times with different labeled sets, and we report the average performance of each model.
Appendix C Comparison Against kNearestNeighbor
@@
Appendix D Experiment on All BODMAS
@@
    The BDOMAS dataset contains many families with only one sample. After a random train/test split, only 296 families are present in both train and test sets. To provide consistent evaluation metrics, we trained and tested on these 296 families only. MalMixer achieves $64.31$, 49.049.049.049.02, $54.28$, and $45.70$ in accuracy, precision, recall, and f1 (macro).
@@
    In another experimental design, we conducted a random train/test split among all 581 malware families, so both sets contained families from outside the top-100. MalMixer was trained using this data, with only $1\%$ of the top-100 prevalent families labeled. During testing, any family not within the top-100 was considered a False Positive if it was classified among those top-100 families. This set-up yielded performance metrics of $69.58$, $60.93$, $60.92$, and $56.11$ on the 100 labeled families. For context, in a prior experiment where both train and test sets solely contained the top-100 families, the performance metrics stood at $70.65$, $63.60$, $62.11$, and $58.78$.
The above results indicate MalMixer handles smaller families reasonably well, and its performance on large families remains strong even with False Positives from small, unidentified families.
@@
TABLE IX: Results grouped by family from the BODMAS-20 experimental design with $1\%$ of labeled data. We note that the few families with lower f1 scores are intrinsically harder to classify, even by the standards of a fully-supervised model.
@@
Appendix E Obfuscation Experiment
Models Accuracy F1 macro
MalMixer 45.64$\pm$0.77 34.97$\pm$0.57
MixMatch 43.31$\pm$1.10 31.35$\pm$1.19
MixUp 42.55$\pm$1.02 28.89$\pm$0.86
VAT 40.68$\pm$0.81 22.47$\pm$2.05
To evaluate MalMixer’s performance under code obfuscation, we used the UCSB-Packed dataset [43], which employs 10 unique packers to create obfuscated malware samples. As the dataset does not have malware family labels, we turned to VirusTotal reports to assign family labels to each sample. We limited our dataset to only the top 20 most common families, and perform a random train-test split. We compare MalMixer against baseline models using only $1\%$ labeled sample in the train set. As shown in Table X, MalMixer can better classify obfuscated sample given few labeled samples than all baselines.
It’s vital to recognize that the performances in Table X only partially capture the increased challenges in classifying obfuscated samples. The noises, inconsistencies, and inaccuracies inherent in VirusTotal’s family labels further complicates matters [25]. Experimentally, we’ve noted that even a fully supervised model struggles to achieve high accuracy on this dataset. This calls for future reserach efforts on curating obfuscated datasets with well-labeled families, as well as on improving static-feature-based classifiers’ robustness to obfuscation.
Appendix F Interpolatable vs. Non-interpolatable feature sets for EMBER features
Interpolatable Features: {ByteHistogram, ByteEntropyHistogram [50], frequency of printable characters, occurrences of strings (e.g. ”HTTP://”, ”MZ”), length of the bytes of the file, virtual size of the life parsed binary, number of exported/imported functions, number of symbols, size of code/headers/heap commits, total number of sections, number of sections with a certain feature (e.g. zero size, empty name), size/virtual size of data directories}

### v4

@@ Appendix A Implementation Details
We use two Nvidia A30 GPUs during model training and evaluation. We set the random seed as 17 in all processes for reproducibility. We note that training and evaluation across all experimental settings required less than one hour each using BODMAS-20 with $1\%$ labeled data. In our experiments, we standardize all features in the BODMAS dataset through z-score scaling as in standard practice. For the encoder-decoder architecture, encoder $\phi_{i}$ comprises three layers, with each layer outputting 512 features. The corresponding decoder $\psi_{i}$ consists of three layers, mirroring the structure of the encoder in reverse order. Encoder $\phi_{n}$ has five layers, outputting 1024, 1024, 512, 512, and 512 features respectively. The corresponding the decoder $\psi_{n}$ reflects the encoder’s architecture in reverse order. Consequently, the embeddings $h_{u}^{d},h_{u}^{s},h_{i}^{d},h_{i}^{s}$ created by the encoder-decoder architecture are all represented as 256-dimensional vectors. For our retrieval-based data augmentation, we use Faiss [60], an open-source library designed for scalable similarity search and clustering, to construct the k-nearest-neighbor graph and perform similarity searches with L2 distance. Top 5 neighbors are selected for the $k$-NN graph in the original feature space $\mathcal{S}$ and 5 candidates are selected for alignment-based malware selection.
For the semi-supervised framework, unless otherwise noted, we use a Fully Connected ResNet (FC-ResNet) [58] as the base classifier for MalMixer and all other semi-supervised methods evaluated. Our FC-Resnet model begins with a fully connected layer projecting the 2,381-dimensional input to a 1,024-dimensional latent space, followed by 12 residual blocks organized into four groups with progressively decreasing dimensions ($1,024-512-256-128$), where each group contains three residual blocks. Each block consists of two linear layers, LeakyReLU activations, and skip connections to aid gradient flow. Downsampling is applied when dimensions change. The final representation is flattened and passed through a fully connected classification head for multi-class prediction. Unlike traditional ResNets in computer vision, our model omits convolutions. We empirically found that removing all BatchNorm layers in FC-ResNet leads to improved performance of MalMixer.
We train MalMixer using an Adam optimizer (learning rate 5e-5 and weight decay 1e-5. For all training and evaluation of MalMixer, we linearly ramp up $\lambda$ to its maximum value [61]. We perform up-sampling through data augmentation to each malware family in the labeled set to create a balanced $\mathcal{X}^{\prime}$. Each evaluation task is performed five times with different labeled sets, and we report the average performance of each model.
Algorithm 1 MalMixer data augmentation pipeline
0:  Malware feature set $\mathcal{S}=concat(\mathcal{S}_{i},\mathcal{S}_{n})$, non-interpolatable feature embedding set $\mathcal{H}_{n}$, feature invariance embedding sets $\mathcal{H}_{i}^{sim}$ and $\mathcal{H}_{n}^{sim}$, pre-trained encoder $\phi_{i}$
1:  for $s\in S$ do
2:     Randomly select a similar malware $s^{\prime}$ from the set of the top-k similar malware for $s$
3:     Linearly combine the interpolatable features of $s$ and $s^{\prime}$ using Eq. (2). Let $\tilde{s}_{i}$ denote the result
4:     Retrieve from $\mathcal{H}_{n}$ the projections of the non-interpolatable features of $s$ and $s^{\prime}$, denote these projections $h_{n}$ and $h_{n}^{\prime}$
5:     Linearly combine $h_{n}$ and $h_{n}^{\prime}$ using Eq. (3) to obtain $\tilde{h}_{n}$
6:     Identify $k~{}candidates$ in $\mathcal{H}_{n}$ most similar to $\tilde{h}_{n}$
7:     for $\hat{h}_{n}$ in $candidates$ do
8:        With $\hat{h}_{n}=\hat{h}_{n}^{sim}\oplus\hat{h}_{n}^{dis}$, set $\hat{h}_{n}=\hat{h}_{n}^{sim}$
9:     end for
10:     Use $\phi_{i}$ to project $\tilde{s}_{i}$ to $\mathcal{H}_{i}$. Let $\tilde{h}_{i}$ denote the result
11:     With $\tilde{h}_{i}=\tilde{h}_{i}^{sim}\oplus\tilde{h}_{i}^{dis}$, set $\tilde{h}_{i}=\tilde{h}_{i}^{sim}$
12:     Select $\hat{h}_{n}$ from $candidates$ s.t. $\hat{h}_{n}$ is most similar to $\tilde{h}_{i}$
13:     Identify the $\tilde{s}_{n}\in\mathcal{S}_{n}$ that maps to $\hat{h}_{n}$. Use $\tilde{s}_{n}$ as the mixture of $s$ and $s^{\prime}$ by their non-interpolatable features.
14:     return  $\mathcal{A}(s):=\tilde{s}_{i}\oplus\tilde{s}_{n}$
15:  end for
TABLE IX: Saturation evaluation of MalMixer against MOTIF-50, with Fully Connected ResNet and LightGBM provided as a supervised standard for reference.
Models Labels Accuracy Precision Recall F1
MalMixer $1\%$ 40.45$\pm$4.77 45.45$\pm$6.59 43.06$\pm$5.31 39.70$\pm$5.16
$5\%$ 46.13$\pm$5.04 47.20$\pm$6.73 45.26$\pm$6.02 42.31$\pm$5.91
$10\%$ 49.33$\pm$5.46 51.52$\pm$3.42 46.71$\pm$3.26 44.71$\pm$4.17
$25\%$ 63.96$\pm$2.78 64.22$\pm$2.77 62.44$\pm$2.664 59.82$\pm$3.00
$50\%$ 71.31$\pm$2.50 71.18$\pm$1.77 70.56$\pm$1.96 68.31$\pm$2.04
$75\%$ 74.18$\pm$1.23 75.81$\pm$2.10 74.08$\pm$1.48 71.59$\pm$1.28
FC-ResNet [58] $100\%$ 81.15 81.78 79.30 77.81
LightGBM [59] $100\%$ 83.07 79.94 79.72 77.40
Appendix B Saturation Analysis on of MOTIF-50
We performed an additional saturation analysis on MOTIF-50, analogous to that on BODMAS-20 in Section 7.1. We select $1\%$, $5\%$, $10\%$, $50\%$, and $75\%$ of all data in training set as labeled data, ensuring that at least one sample from each family is labeled. In the $1\%$ setting, almost all families contain only one-shot labels. Table IX shows that achieving satisfactory performance on MOTIF-50 requires a higher proportion of labeled data compared to BODMAS-20 (Table IV). However, we note that even with 25% labeled data, most families in MOTIF-50 still remain in the few-shot regime (i.e. fewer than 5 labels per family) due to the larger number of families and the small family sizes in this dataset.
More importantly, we hypothesize that MalMixer’s degrading performance with fewer labels on MOTIF-50 may be more reflective of label quality issues rather than the true impact of the reduced label quantity. Specifically, MOTIF’s label noise is reported to be as high as 40% [47]. As the number of labeled samples decreases, the model become more reliant on the quality of each label; in contrast, with larger labeled sets, the distribution of labeled samples within each family naturally converges to the true mean, even in the presence of high label noises. Thus, in few-shot settings, it is typically assumed that labels are highly accurate, as experts have time to carefully curate them. These factors suggest that MOTIF-50 may not be the ideal dataset for evaluating the impact of decreasing labeled data quantities on model performance.
Nonetheless, even under such settings, MalMixer still achieves satisfactory performance under few-shot scenarios and outperforms baselines in near one-shot settings under high noise, as shown Section 7.2.
Appendix C Expressiveness of FC-ResNet
Fully Connected ResNet (FC-ResNet), our chosen base classifier, achieved performance comparable to LightGBM in fully supervised settings on both BODMAS-20 and MOTIF-50, as shown in Table IV and Table IX. In tabular machine learning, gradient boosting techniques like LightGBM are often considered state-of-the-art due to their strong performance. In fact, LightGBM was the primary classification method in the original BODMAS paper [24]. However, our semi-supervised framework requires a base classifier that outputs label probabilities, making GBDT (e.g. LightGBM) unsuitable for this role.
To explore alternative expressive models, we experimented with a Transformer-based classifier designed specifically for tabular malware detection. This model projects input features to a 128-dimensional latent space, applies positional encoding, and utilizes two Transformer encoder blocks with four attention heads and a hidden dimension of 256 in the feedforward layers. However, despite its theoretical expressiveness, the Transformer achieved only 91.16 Accuracy, 91.89 Precision, 92.80 Recall, and 92.22 F1 under fully supervised settings on BODMAS-20, underperforming compared to FC-ResNet (Table IV). This is possibly due to limtied dataset sizes and suggests that a more expressive neural model does not necessarily translate to better performance for this tabular malware classification task.
As further evidence, prior studies have demonstrated the effectiveness of a simple three-layer MLP as a base classifier on similar tasks [47]. Collectively, these findings confirm that FC-ResNet offers sufficient expressiveness to serve as the base classifier in our semi-supervised framework.
Appendix D Comparison Against kNearestNeighbor
@@
Appendix E Experiment on All BODMAS
@@
    Experiment 1 (All Families Labeled). The BODMAS dataset contains many families with only one sample. After a random 80/20 train/test split, only 296 families are present in both train and test sets. To provide consistent evaluation metrics, we trained and tested on these 296 families only. We labeled $1\%$ of the samples, ensuring at least one labeled sample per family MalMixer achieves $64.31$, 49.049.049.049.02, $54.28$, and $45.70$ in accuracy, precision, recall, and f1 (macro). For reference, the fully supervised FC-ResNet yielded performance metrics $86.93$, $63.29$, $61.05$, $60.36$. While MalMixer’s performance drops substantially in the presence of numerous small families, this outcome is unsurprising given the challenge of handling more small families, especially when benchmarked against the also deteriorated performance of a fully supervised approach.
@@
    Experiment 2 (All Families, Only Top-100 Families Labeled). Here, we conducted a 80/20 random train/test split among all 581 malware families, so both sets contained families from outside the top-100 in prevalence. MalMixer was trained using this data, with only $1\%$ of the top-100 prevalent families labeled. During testing, any family not within the top-100 was considered a False Positive if it was classified among those top-100 families. This set-up yielded performance metrics of $69.58$, $60.93$, $60.92$, and $56.11$ on the 100 labeled families. For context, in a prior experiment where both train and test sets solely contained the top-100 families, the performance metrics stood at $70.65$, $63.60$, $62.11$, and $58.78$. These results suggest that MalMixer’s performance on large families remains strong even with False Positives from small, unidentified families. In general, the presence of small families does not significantly impact the model’s performance on large families. We hypothesize that the minimal size of these long-tail families limits their overall impact on both training and inference metrics.
@@
Appendix F Obfuscation Experiment
To evaluate MalMixer’s performance under code obfuscation, we used the UCSB-Packed dataset [43], which employs 10 unique packers to create obfuscated malware samples. As the dataset does not have malware family labels, we turned to VirusTotal reports to assign family labels to each sample. We limited our dataset to only the top 20 most common families, and perform a random 80/20 train-test split. We compare MalMixer against baseline models using only $1\%$ labeled sample in the train set.
Under packing, many of the $k$-NN and clustering assumptions outlined in Sections 1 and 3.3 become weaker, negatively impacting MalMixer’s performance on the UCSB-Packed dataset. However, as shown in Table X, MalMixer remains among the top two methods, performing slightly worse in accuracy and precision but better in recall and F1 compared to MixMatch (Gaussian). Notably, MixMatch (Gaussian) is structurally identical to MalMixer but replaces our proposed data augmentation with Gaussian noise augmentation. This suggests that even when some of MalMixer’s underlying assumptions are violated, our proposed data augmentation strategy remains at least competitive with Gaussian-noise-based augmentation, which does not respect malware feature semantics.
It’s vital to recognize that the performances in Table X only partially capture the increased challenges in classifying obfuscated samples. The noises, inconsistencies, and inaccuracies inherent in VirusTotal’s family labels further complicates matters [24]. Experimentally, we’ve noted that even a fully supervised FC-ResNet achieved only 72.46 accuracy on this dataset. This calls for future research efforts on curating obfuscated datasets with less noisy family labels, as well as on improving static-feature-based, few-shot, and semi-supervised classifiers’ robustness to obfuscation.
Models Accuracy Precision Recall F1
VAT 40.68$\pm$0.81 32.03$\pm$3.67 23.06$\pm$1.61 22.47$\pm$2.05
VAT (aug) 38.89$\pm$0.86 32.13$\pm$1.13 31.93$\pm$1.61 29.60$\pm$0.35
MixUp 42.55$\pm$1.02 35.53$\pm$2.14 29.33$\pm$1.20 28.89$\pm$0.86
MixMatch 43.31$\pm$1.10 37.03$\pm$1.02 31.28$\pm$1.37 31.35$\pm$1.19
MixMatch (Gaussian) 46.80$\pm$0.73 39.43$\pm$2.16 34.84$\pm$0.97 34.07$\pm$0.25
MORSE 43.59$\pm$2.31 42.88$\pm$2.84 27.77$\pm$2.38 28.62$\pm$2.89
MalMixer 46.08$\pm$0.76 37.41$\pm$0.74 35.70$\pm$0.69 35.24$\pm$0.53
Appendix G Interpolatable vs. Non-interpolatable feature sets for EMBER features
Interpolatable Features: {ByteHistogram, ByteEntropyHistogram [50], frequency of printable characters, occurrences of strings (e.g. “HTTP:/”, “MZ”), length of the bytes of the file, virtual size of the LIEF-parsed binary, number of exported/imported functions, number of symbols, size of code/headers/heap commits, total number of sections, number of sections with a certain feature (e.g. zero size, empty name), size/virtual size of data directories}
Appendix H Train-Test Overlap of Non-Interpolatable Features in Experimental Datasets
One important detail about our experimental dataset is the train-test overlap of non-interpolatable features. This information could clarify whether the models are relying too much on non-interpolatable features..
For BODMAS-20, 81% of non-interpolatable features within the same malware family overlap between the training and test sets. In contrast, the same number for MOTIF-50 is just 6%. This partially explains why all models, including MalMixer and baselines, performed better on BODMAS-20. However, MalMixer outperforms other models on both datasets, showing its ability to handle both high and low train-test overlaps in non-interpolatable features. This reasonably suggests that our model does not overly rely on non-interpolatable features, at least compared to the other baseline models.
Furthermore, 37% of non-interpolatable features are unique per malware family in BODMAS-20. The same number for MOTIF-50 is 97%. These numbers suggest that our data augmentation pipeline for non-interpolatable features can indeed replace one sample’s non-interpolatable features with other distinct non-interpolatable feature sets, thus introducing diversity to mitigate overfitting and increase robustness.
Algorithm 2 The modified MORSE algorithm.
0:  Training dataset $D$, number of total training epochs $K$, labeled data’s proportion $d$, starting re-weighting epoch $T_{d}$, confidence threshold $\tau$, unsupervised loss weight $\lambda$, learning rate $\alpha$.
1:  Initialization: Initialize the weights $\Theta$ for the model $f(\cdot)$ by using the entire dataset $D$ with only a few epochs.
2:  for $k=0,1,2,\dots,K$ do
3:     Partition the training dataset $D$ into labeled dataset $\mathcal{X}$ and unlabeled dataset $\mathcal{U}$: select the top $d\%$ examples with the least loss values from each given class and treat them as labeled data, and the rest as unlabeled data.Use the groundtruth labeled dataset $\mathcal{X}$ and unlabeled dataset $\mathcal{U}$.
4:     for $iter=1,2,\dots,num\_batches$ do
5:        From $\mathcal{X}$, draw a mini-batch:$\{(\mathbf{x}_{b},y_{b}):b\in(1,\dots,B)\}$
6:        From $\mathcal{U}$, draw a mini-batch:$\{\mathbf{u}_{b}:b\in(1,\dots,B_{\mu})\}$
7:        for $b=1,2,\dots,B_{\mu}$ do
8:           // Perform weak augmentation against $\mathbf{u}_{b}$
9:           $\mathbf{q}_{b}=f(g(\mathbf{u}_{b});\Theta)$
10:           $\widehat{q}_{b}=\arg\max(\mathbf{q}_{b})$
11:        end for
12:        if $k<T_{d}$ True then
13:           // Perform weak augmentation against $\mathbf{x}_{b}$
14:           $\mathcal{L}_{s}=\frac{1}{B}\sum_{b=1}^{B}H(y_{b},f(g(\mathbf{x}_{b})))$
15:           // Perform strong augmentation against $\mathbf{u}_{b}$
16:           $\mathcal{L}_{u}=\frac{1}{B_{\mu}}\sum_{b=1}^{B_{\mu}}\textbf{I}(\max(\mathbf{q}_{b})>\tau)H(\widehat{q}_{b},f(h(\mathbf{u}_{b})))$
17:        else
18:           // Calculate the weight of each training sample
19:           // Perform weak augmentation against $\mathbf{x}_{b}$
20:           $\mathcal{L}_{s}=\frac{1}{B}\sum_{b=1}^{B}w_{b}H(y_{b},f(g(\mathbf{x}_{b})))$
21:           // Perform strong augmentation against $\mathbf{u}_{b}$
22:           $\mathcal{L}_{u1}=\frac{1}{B_{\mu}}\sum_{b=1}^{B_{\mu}}\textbf{I}(\max(\mathbf{q}_{b})>\tau)$
23:           $\mathcal{L}_{u}=\mathcal{L}_{u1}+w_{b}H(\widehat{q}_{b},f(h(\mathbf{u}_{b})))$
24:        end if
25:        $\mathcal{L}=\mathcal{L}_{s}+\lambda\mathcal{L}_{u}$
26:        Update the model’s weights $\Theta$ by minimizing the loss function $\mathcal{L}$
27:        Optional: decay the learning rate $\alpha$
28:     end for
29:  end for
30:  Output: the well-trained model $f(\cdot;\Theta)$.
Appendix I Adaptations of MORSE
MORSE (Malware cOrpus foR Semi-supervised lEarning) [47] is a state-of-the-art malware family classification algorithm, originally designed to address challenges with incorrect groundtruth malware labels and class imbalance. Leveraging semi-supervised learning, MORSE gracefully handles malware classifications in label-scarce scenarios using both labeled and unlabeled data.
At a high level, MORSE’s original design starts with a fully labeled dataset but retains only the most confidently classified samples as labeled data, treating the rest as unlabeled. Then, the labeled and unlabeled data is fed into the FixMatch [62] pipeline, a semi-supervised learning approach originally developed for image classification and a successor to MixMatch [23]. Within this pipeline, MORSE introduces sample re-weighting [63] to mitigate class imbalance. Additionally, it employs a malware-specific data augmentation technique: an augmented sample is obtained by swapping a subset of randomly selected features from a given sample with those from another randomly selected sample. To align with FixMatch’s requirement for both weak and strong augmentations, MORSE defines weak augmentation as replacing $10\%$ of features and strong augmentation as replacing $20\%$. Note that although MORSE proposes a malware-specific data augmentation pipeline, the random swapping still largely disregards the underlying semantics of each feature.
Unlike MalMixer, MORSE utilizes a 3-layer MLP as its base classifier. We chose not to replace it with our FC-ResNet, as our experiments showed no performance gains from this substitution.
To adapt MORSE to our experimental settings, we note that our few-shot settings, with already scarce labels, do not permit the model to decide what data should be considered labeled vs. unlabeled. Furthermore, the few-shot settings have rendered the re-sampling techniques empirically unhelpful, partially due to difficulties with inferring class sizes based on few-shot labels. As a result, we have removed these features from MORSE, and have detailed our changes to the original MORSE framework in Algorithm 2, where steps removed are marked with strikethrough and added steps highlighted in gray.
With careful fine-tuning, we found that strictly following the hyperparameter set used in MORSE’s original experiments — which were conducted on their synthetic dataset curated from a subset of BODMAS — while only adjusting the confidence threshold $\tau$ to $0.95$, yielded the best results in our experimental setting.
@@ Appendix I Adaptations of MORSE
TABLE XI: Results grouped by family from the BODMAS-20 experimental design with $1\%$ of labeled data. We note that the few families with lower f1 scores are intrinsically harder to classify, even by the standards of a fully-supervised model.